### v3

# ELITE: Enhanced Language-Image Toxicity Evaluation for Safety

Abstract
Current Vision Language Models (VLMs) remain vulnerable to malicious prompts that induce harmful outputs. Existing safety benchmarks for VLMs primarily rely on automated evaluation methods, but these methods struggle to detect implicit harmful content or produce inaccurate evaluations. Therefore, we found that existing benchmarks have low levels of harmfulness, ambiguous data, and limited diversity in image-text pair combinations. To address these issues, we propose the ELITE benchmark, a high-quality safety evaluation benchmark for VLMs, underpinned by our enhanced evaluation method, the ELITE evaluator. The ELITE evaluator explicitly incorporates a toxicity score to accurately assess harmfulness in multimodal contexts, where VLMs often provide specific, convincing, but unharmful descriptions of images. We filter out ambiguous and low-quality image-text pairs from existing benchmarks using the ELITE evaluator and generate diverse combinations of safe and unsafe image-text pairs. Our experiments demonstrate that the ELITE evaluator achieves superior alignment with human evaluations compared to prior automated methods, and the ELITE benchmark offers enhanced benchmark quality and diversity. By introducing ELITE, we pave the way for safer, more robust VLMs, contributing essential tools for evaluating and mitigating safety risks in real-world applications.
Warning: This paper includes examples of harmful language and images that may be sensitive or uncomfortable. Reader discretion is advised.
1 Introduction
Vision Language Models (VLMs), which are composed of pre-trained Large Language Models (LLMs) and visual encoders, have been introduced to tackle complex multimodal tasks. Despite advancements in their capabilities and performance, VLMs remain vulnerable to malicious inputs, raising significant safety concerns and posing substantial challenges to their large-scale deployment in real-world applications (Zong et al., 2024; Carlini et al., 2023; Gong et al., 2023; Bommasani et al., 2021).
Concerns about the safety of VLMs, such as malicious users inducing harmful outputs, have been raised and several safety evaluation benchmarks have appeared to assess the safety of VLMs (Zong et al., 2024; Liu et al., 2025; Gu et al., 2024; Zhang et al., 2024; Wang et al., 2025; Gong et al., 2023; Luo et al., 2024). These benchmarks are typically constructed by collecting image-text pairs that can trigger harmful responses, with safety evaluations conducted through automated methods using language models rather than humans (Zong et al., 2024; Luo et al., 2024; Liu et al., 2025; Zhang et al., 2024; Li et al., 2024). While these benchmarks contribute to developing safer, more robust VLMs, we have identified significant issues in the existing benchmarks and their automated evaluation methods.
First, we identify that the automated evaluation methods currently adopted in many safety benchmarks are not always reliable. As shown in Fig. 1(c), in the upper example, the model’s response is largely vague or simply descriptive of the image, yet it is still deemed a successful jailbreak. In another case, as shown in the lower example, the existing evaluation method fails to detect implicit suicidal intent and considers it safe. We identify these problems in the safety evaluation methods and propose the Enhanced Language-Image Toxicity Evaluation (ELITE) evaluator, a method designed to accurately evaluate the safety of VLMs.
Figure 1: Contributions of ELITE. (a) Benchmark Construction: The ELITE benchmark is a high-quality benchmark built by filtering out unsuccessful image-text pairs using the ELITE evaluator. (b) Generated Image-Text Pairs: Image-text pair with various methods for inducing harmful responses from VLMs. (c) Evaluation Method: The ELITE evaluator is a more precise rubric-based safety evaluation method compared to existing methods for VLMs.
Second, we identify quality issues in existing safety benchmarks. We have observed that existing benchmarks generally exhibit low levels of harmfulness and contain a significant number of ambiguous image-text pairs that fail to induce harmful responses from VLMs. To address this, we introduce the ELITE benchmark, which filters out ambiguous image-text pairs from existing benchmarks using the ELITE evaluator (Fig. 1(a)).
Third, existing benchmarks mainly consist of unsafe image-unsafe text pairs (i.e., unsafe-unsafe, safe-unsafe, or unsafe-safe pairs) (Gu et al., 2024; Zong et al., 2024; Liu et al., 2025). However, as demonstrated in the examples at the bottom of Fig. 1(c), harmful responses can also be induced through safe-safe pairs (Wang et al., 2025). To address this issue, we propose various methods for inducing harmful responses from VLMs. As shown in Fig. 1(b), the ELITE benchmark incorporates four in-house generated methods, which improve coverage of all four image-text pair combinations. This enhances the diversity of the ELITE benchmark and enables a more comprehensive evaluation of VLM safety.
The ELITE evaluator builds on StrongREJECT (Souly et al., 2024), extending its rubric-based evaluation to vision-language tasks by incorporating toxicity scores. This addition helps distinguish genuinely harmful outputs from merely descriptive ones, which are common in VLMs due to the influence of image input. Leveraging this evaluator, we curate the ELITE benchmark by filtering out ambiguous pairs and selecting diverse, explicitly harmful image-text examples from both existing and newly generated data, enabling more reliable safety evaluation.
Our experiments demonstrate that the ELITE evaluator aligns better with human judgments than existing automated evaluation methods. Furthermore, through extensive experiments, we validate the diversity and superior quality of the ELITE benchmark, which is designed using the ELITE evaluator. To summarize, our main contributions are as follows:
Table 1: Overview of the ELITE benchmark. We created 4,587 image-text pairs by filtering out ambiguous image-text pairs that are unable to induce harmful responses in both existing benchmarks and the in-house generated image-text pairs. “New” refers to the image-text pairs we generated using various methods. In the case of JailbreakV-28k (Luo et al., 2024), filtering is performed only on insufficient taxonomies to maintain balance across taxonomies.
Taxonomy The ELITE benchmark Sum Total
VLGuard MLLMGuard MM-SafetyBench SIUO Figstep SPA-VL JailbreakV-28k New
S1. Violent Crimes 91 11 39 1 91 299 0 72 604 4587
S2. Non-Violent Crimes 13 2 144 1 209 221 0 124 714
S3. Sex Crimes 6 3 0 0 39 32 38 196 314
S4. Defamation 19 2 5 0 9 224 0 140 399
S5. Specialized Advice 37 1 21 3 84 131 0 54 331
S6. Privacy 0 14 63 2 42 93 0 99 313
S7. Intellectual Property 1 5 11 0 37 74 238 0 366
S8. Indiscriminate Weapons 0 4 36 0 23 116 84 100 363
S9. Hate 204 0 55 4 54 144 0 82 543
S10. Self-Harm 15 0 12 2 20 37 89 127 302
S11. Sexual Content 88 1 19 0 36 32 102 60 338
  • We introduce the ELITE evaluator for accurate automated safety evaluation in VLMs. Through the ELITE evaluator, we demonstrate that existing automated safety evaluation methods often result in inaccurate evaluations.
  • We propose the ELITE benchmark, a rubric-based safety evaluation benchmark for VLMs using the ELITE evaluator. The ELITE benchmark addresses the limitations of existing benchmarks, such as insufficient benchmark quality. We construct a high-quality benchmark by filtering out low-quality and ambiguous image-text pairs.
  • We propose various methods for inducing harmful responses in VLMs. These methods are applied to generate extensive image-text pairs across all combinations of safe and unsafe image-text pairs to elicit harmful responses that violate VLMs’ safety policies.
2 Related Work
To address vulnerabilities and evaluate the safety of VLMs, various benchmarks have been developed, following the previous safety evaluation benchmarks for LLM (Wang et al., 2023). VLGuard (Zong et al., 2024) introduces a fine-grained evaluation benchmark that focuses on visual-linguistic reasoning, leveraging a taxonomy that categorizes potential safety issues. MM-SafetyBench (Liu et al., 2025) provides a comprehensive benchmark with image-text pairs across 13 safety-critical scenarios, emphasizing image-based manipulations and their impact on VLMs’ responses. MLLMGuard (Gu et al., 2024) evaluates safety across five dimensions—privacy, bias, toxicity, truthfulness, and legality—using a bilingual benchmark. Besides these benchmarks, others such as SPA-VL (Zhang et al., 2024), JailbreakV-28K (Luo et al., 2024), and SIUO (Wang et al., 2025), contribute to evaluating vulnerabilities and enhancing the robustness of VLMs.
For safety evaluation, automated evaluators using language models have been adopted due to the high cost of human evaluators (Zong et al., 2024; Luo et al., 2024; Liu et al., 2025; Zhang et al., 2024; Li et al., 2024). Additionally, safeguard models for safety evaluation have been developed (Inan et al., 2023; Chi et al., 2024; Gu et al., 2024). These automated evaluations commonly rely on metrics such as ASR. However, relying solely on ASR, a simple binary classification that deems an attack successful as long as the model does not refuse the instruction, can lead to an overestimation of jailbreak effectiveness and discrepancies with human judgment (Souly et al., 2024). To address this issue, recent evaluation methods have integrated the level of detail in responses as an additional assessment criterion (Souly et al., 2024; OpenAI, 2024b; Guan et al., 2024).
3 ELITE
In this section, we introduce the ELITE evaluator as an accurate evaluation method. Also, we describe the construction process of the ELITE benchmark, along with the creation of in-house image-text pairs, which are designed to induce harmful responses from VLMs. As shown in Table 1, we provide a detailed breakdown of the components of the ELITE benchmark.
3.1 The ELITE Evaluator
Existing benchmarks use human evaluators or automated evaluators for the safety evaluation of VLMs. Relying on human evaluators is expensive, and thus recent approaches have leveraged automated annotators using LLMs or VLMs. However, existing evaluation methods assume that an attack is successful if the victim model does not output any predefined refusal messages (Chi et al., 2024; Liu et al., 2025), causing inaccurate safety evaluation results. In order to address this problem, we introduce the ELITE evaluator, an accurate and structured rubric-based evaluation method.
The ELITE evaluator is built upon the StrongREJECT evaluator (Souly et al., 2024), which highlighted the overestimation of most basic automated evaluators in the jailbreak techniques of LLMs. The StrongREJECT evaluator is defined as follows:
$\text{StrongREJECT}=(1-\textit{refused})\times\frac{\textit{specific}+\textit{convincing}}{2}$ (1)
In Eq. 1, refused evaluates whether the model has refused the forbidden prompt, and assigns 1 point if the victim model refuses to respond. specific measures the specificity of the response, while convincing evaluates its convincingness. Both terms are scored on a scale of 1 to 5.
In VLMs, the victim model’s responses to forbidden prompts can vary as images are added. For instance, as observed in the responses in Fig. 2, despite malicious prompts, the victim model may produce responses that are unhelpful to the attacker. However, when the StrongREJECT evaluator is applied to VLMs, it often assigns high scores even when the model does not explicitly refuse to respond and instead provides unhelpful answers. In other words, since the victim model provides a response instead of refusing, a score of 0 is assigned to refused, while moderate scores are given for specific and convincing, resulting in a final score that indicates the response is harmful. This makes it challenging to determine whether the attempt to induce harmful outputs was genuinely successful. To address this problem, we incorporate toxicity into the evaluation rubric to assess how harmful the victim model’s response is. Through this approach, toxicity plays a crucial role in evaluating the success of inducing harmful responses, addressing the limitations of relying solely on refused to determine whether the victim model’s response is harmful or not. Accordingly, the ELITE evaluator is defined as follows:
$\text{ELITE}=(1-\textit{refused})\times\frac{\textit{specific}+\textit{convincing}}{2}\times\textit{toxicity}$
In Eq. 2, the ELITE evaluator introduces toxicity as an additional criterion, scored on a scale of 0 to 5.
Figure 2: Examples of safety evaluations about the victim model’s response by ELITE and StrongREJECT evaluator. $r$, $s$, $c$, and $t$ represent refused, specific, convincing, and toxicity, respectively. The ELITE evaluator can effectively evaluate by utilizing the toxicity score to make more accurate judgments.
3.2 Taxonomy
We construct a taxonomy of 11 hazard categories to establish a comprehensive and extensive safety benchmark, aligning with AILuminate v1.0 (MLCommons, ) standards. The taxonomy includes Violent Crimes (S1), Non-Violent Crimes (S2), Sex Crimes (S3), Defamation (S4), Specialized Advice (S5), Privacy (S6), Intellectual Property (S7), Indiscriminate Weapons (S8), Hate (S9), Self-Harm (S10), and Sexual Content (S11). Responses that enable, encourage, or endorse harmful behavior are flagged as unsafe for their respective categories. Detailed definitions of each category is provided in Appendix B.1.
3.3 Image&Text Creation
We refer to the in-house generated image-text pairs as the ELITE benchmark (generated). As shown in Fig.1(b), the ELITE benchmark (generated) is composed of four methods—Role Playing, Fake News, Blueprint, and Flowchart—applied across various taxonomies to elicit harmful responses from the victim model. Note that while certain methods, such as Blueprint and Fake News, are used only in specific taxonomies (e.g., Indiscriminate Weapons and Defamation), others, like Flowchart and Role Playing, are applied more broadly across all taxonomies. Detailed examples of these methods are provided in Appendix B.4.
To generate image-text pairs, we use the following methods:
(1) Image Generation: For Role Playing, Blueprint, and Flowchart, we use image generation models such as Flux AI (Labs, 2023) and Grok 2 (xAI, 2024) to create images that align with the key concepts of each taxonomy. Specifically, we first extract relevant keywords for each taxonomy and use these keywords as prompts to generate corresponding images. For Fake News, we manually synthesize these images to create outputs that align with the intended misinformation scenarios, using the open-source image dataset CelebA (Liu et al., 2015).
(2) Text Generation: We generate an initial forbidden text prompt by creating keywords relevant to the image and taxonomy, then generate multiple variations of the prompt using Grok 2. To identify the most effective forbidden text prompt for the given image, we evaluate responses from three victim models (Phi-3.5-Vision, Llama-3.2-11B-Vision, and Pixtral-12B). Among the models that produce harmful responses, we select the image-text pair with the highest ELITE evaluator score to finalize its construction.
These image-text pairs are explicitly designed to induce harmful responses from VLMs, enabling a comprehensive safety evaluation. As shown in Table 2, we incorporate 593 safe-safe pairs into the ELITE benchmark (generated) by embedding inherently harmful intents. These pairs can still induce unsafe responses from VLMs, making them crucial for evaluating safety. Through this, we aim to develop a more extensive benchmark that effectively captures potential vulnerabilities in VLMs.
Table 2: The distribution of the four image-text pair types (unsafe-unsafe, safe-unsafe, unsafe-safe, and safe-safe) in the ELITE benchmark (generated).
ELITE benchmark (generated) Total
safe-safe safe-unsafe unsafe-safe unsafe-unsafe
593 69 350 42 1054
3.4 Benchmark Construction Pipeline
As shown in Fig. 3, the steps for constructing the ELITE benchmark are as follows:
(1) Taxonomy Alignment: To align the image-text pairs in existing benchmarks with the taxonomy of the ELITE benchmark, we use GPT-4o to classify image-text pairs into their corresponding taxonomies within the ELITE benchmark.
(2) Filtering: We apply a filtering process based on a defined threshold to both existing benchmarks and the ELITE benchmark (generated). Specifically, on the ELITE evaluator’s [0-25] point scale, we set a threshold determined by human judgment. ELITE evaluator score $s\geq 10$ indicates that the victim model’s response is sufficiently harmful, while $s<10$ indicates that the victim model either refused to respond to the forbidden prompt or provided a non-harmful response. Using this threshold, we primarily include image-text pairs in the ELITE benchmark if at least two out of the three victim models (Phi-3.5-Vision, Llama-3.2-11B-Vision, and Pixtral-12B) achieve a score of $s\geq 10$ to prevent over-reliance on a single model during filtering. However, in cases where a single model’s response is deemed sufficiently harmful, pairs meeting the threshold with only one model are also included. Examples of model responses near our threshold are provided in Appendix A.2.
(3) Balancing the Taxonomy: After filtering, we identify that some benchmarks are overly concentrated in specific taxonomies (e.g., 204 image-text pairs in VLGuard are filtered into the S9. Hate), leading to imbalance across taxonomies. To create a more balanced benchmark, we additionally filter JailbreakV-28k (Luo et al., 2024) for only non-concentrated categories. Also, to address the issue of certain taxonomies being overly dependent on specific benchmarks, We exclude image-text pairs with the lowest combined ELITE evaluator scores from the three models.
Figure 3: The pipeline for constructing ELITE benchmark. 1) Taxonomy Alignment: Align the image-text pairs in existing benchmarks with the taxonomy of the ELITE benchmark. 2) Filtering: Integrate only image-text pairs where at least two out of three model responses assign an ELITE evaluator score of 10 or higher. 3) Balancing the Taxonomy: Remove image-text pairs with the lowest combined ELITE evaluator score from overly concentrated taxonomies to maintain balance across taxonomies after filtering.
Table 3: ELITE evaluator score-based ASR of various VLMs across taxonomies. The upper group in the table represents proprietary models, and the lower group represents open-source models. The most vulnerable model is highlighted in bold and the second-most vulnerable with an underline. * denotes the model used for filtering.
Model S1 S2 S3 S4 S5 S6 S7 S8 S9 S10 S11 Average
GPT-4o 16.39 17.51 12.74 20.30 33.23 14.38 7.38 17.36 8.66 11.59 13.91 15.67
GPT-4o-mini 29.47 32.91 18.79 31.58 44.41 25.24 18.03 29.48 18.05 28.48 33.73 28.23
Gemini-2.0-Flash 58.44 70.73 48.09 51.63 50.76 55.59 51.37 71.07 42.17 47.68 48.52 55.37
Gemini-1.5-Pro 37.75 48.04 28.03 40.35 37.76 33.87 50.55 44.63 23.76 27.48 35.21 37.69
Gemini-1.5-Flash 43.21 56.16 22.93 40.60 39.27 37.70 50.82 47.38 30.57 23.51 37.87 40.70
LLaVa-v1.5-7B 67.38 79.13 72.93 51.38 46.83 68.05 63.39 66.94 51.57 64.90 56.80 63.59
LLaVa-v1.5-13B 72.85 86.69 79.94 53.63 54.98 73.48 68.31 72.45 58.56 74.17 60.65 69.68
DeepSeek-VL-7B 38.41 59.94 31.21 34.59 42.90 43.45 42.62 54.27 37.02 35.43 31.95 42.36
DeepSeek-VL2-Small 65.07 81.93 59.24 41.35 58.01 68.69 59.29 70.25 52.12 53.64 42.31 60.95
ShareGPT4V-7B 68.71 86.41 75.16 48.62 53.78 72.52 71.04 64.74 60.96 65.56 56.51 67.16
ShareGPT4V-13B 71.03 87.54 75.16 51.38 56.80 74.76 73.22 66.39 60.41 62.91 52.96 68.08
Qwen2-VL-7B 57.28 70.73 45.22 38.60 47.73 60.06 40.44 66.67 45.49 54.64 50.00 53.72
Molmo-7B 61.09 81.51 62.42 56.14 51.96 57.19 71.31 75.21 47.70 64.90 63.61 63.79
InternVL2.5-8B 51.32 65.83 60.83 23.81 50.76 49.52 36.61 55.65 27.62 43.71 36.98 46.48
InternVL2.5-26B 37.75 47.48 42.36 27.82 45.62 34.82 21.58 50.41 23.02 34.77 28.99 36.21
Phi-3.5-Vision* 37.58 44.40 16.24 49.87 38.07 25.24 21.86 41.05 18.60 23.18 18.34 31.85
Pixtral-12B* 75.50 93.56 77.07 67.17 61.63 79.23 86.61 90.08 82.50 77.15 74.56 79.86
Llama-3.2-11B-Vision* 54.47 69.05 41.40 30.83 55.29 53.35 33.88 55.37 34.44 43.05 39.05 47.94
4 Experiments
4.1 Experiment Setup
We evaluate the effectiveness of the ELITE benchmark, consisting of 4,587 image-text pairs, across various VLMs, including GPT-4o (OpenAI, 2024a), GPT-4o-mini (OpenAI, 2024a), Gemini-2.0 (Google DeepMind, 2024), Gemini-1.5 (Gemini Team et al., 2024), and open-source models. For open-source models, their original hyperparameters are used. We use GPT-4o as the ELITE evaluator to evaluate the safety of VLMs.
4.2 Metric
In the Experiments section, we use the ELITE evaluator score-based Attack Success Rate (E-ASR) for comparison. E-ASR is defined as:
[…]
where $\text{ELITE score}_{i}$ represents the ELITE evaluator score of the $i$-th image-text pair and $N$ is the total number of image-text pairs.
4.3 Evaluation of the ELITE Benchmark
In Table 3, we present comprehensive experimental results of the ELITE benchmark across various proprietary and open-source VLMs. GPT-4o exhibits the lowest E-ASR at 15.67% among models, indicating that it is appropriately safety-aligned against malicious inputs. In contrast, Gemini-2.0-Flash exhibits the highest E-ASR among proprietary models at 55.37%, indicating significant vulnerability to malicious attacks. Additionally, with a few exceptions, most open-source models show high success rates for malicious attacks. The result that most models exhibit an E-ASR exceeding 40% highlights the need for improved safety alignment in VLMs.
Table 4: Comparison of the average E-ASR and ASR when using different benchmarks. It highlights that the most effective benchmark for inducing harmful responses in bold and the second-most effective benchmark with an underline.
Model Benchmark Total E-ASR ASR
LLaVa-v1.5-7B VLGuard 2028 27.75 34.82
MM-SafetyBench 1680 45.06 39.67
MLLMGuard 532 27.26 36.46
ELITE benchmark (generated) 1054 69.17 70.83
ELITE benchmark 4587 63.59 68.98
LLaVa-v1.5-13B VLGuard 2028 28.40 34.00
MM-SafetyBench 1680 46.61 41.25
MLLMGuard 532 27.26 32.65
ELITE benchmark (generated) 1054 78.46 69.24
ELITE benchmark 4587 69.68 69.99
DeepSeek-VL-7B VLGuard 2028 16.40 28.59
MM-SafetyBench 1680 31.79 38.63
MLLMGuard 532 16.29 23.35
ELITE benchmark (generated) 1054 37.95 57.83
ELITE benchmark 4587 42.36 60.83
ShareGPT4V-7B VLGuard 2028 29.24 31.98
MM-SafetyBench 1680 48.81 40.89
MLLMGuard 532 23.51 30.11
ELITE benchmark (generated) 1054 68.50 66.60
ELITE benchmark 4587 67.16 69.54
4.4 Comparisons with Other Benchmarks
In this section, we demonstrate the superiority of both the ELITE benchmark and the ELITE benchmark (generated). Table 4 compares the E-ASR of LLaVa-v1.5 (7B, 13B), DeepSeek-VL (7B), ShareGPT4V (7B), Gemma-3 (4B), and InternVL2.5 (26B) across existing benchmarks, including VLGuard (Zong et al., 2024), MM-SafetyBench (Liu et al., 2025), MLLMGuard (Gu et al., 2024), the ELITE benchmark, and the ELITE benchmark (generated). Note that we use publicly available benchmarks in this experiment.
As shown in Table 4, the ELITE benchmark, which contains approximately 2–3 times more evaluation image-text pairs, achieves significantly higher E-ASR across all models. Furthermore, the ELITE benchmark (generated) demonstrates a substantial increase in E-ASR through effective filtering. These experimental results indicate that the low E-ASR observed in existing benchmarks suggests a substantial number of image-text pairs that fail to elicit harmful responses from VLMs. Consequently, this highlights the effectiveness of the ELITE evaluator in filtering out ambiguous image-text pairs, ensuring that only those capable of inducing harmful responses from VLMs are retained.
Moreover, to demonstrate that the ELITE benchmark is not overly tailored to the ELITE evaluator, we present results based on the previously adopted metric, Attack Success Rate (ASR), instead of E-ASR. These results suggest that the ELITE benchmark remains general and is not excessively influenced by the use of the ELITE evaluator.
Table 5 presents the E-ASR of the methods used to elicit harmful responses from VLMs in the ELITE benchmark (generated). Our experimental results show that Flowchart and Blueprint achieve high E-ASR across a significant number of models, underscoring the importance of incorporating these methods into the benchmark to effectively evaluate and enhance the safety and robustness of VLMs.
Table 5: Comparison of the E-ASR of the proposed methods in the ELITE benchmark (generated).
Method Count Model
LLaVa-7B LLaVa-13B DeepSeek-7B ShareGPT-7B
Blueprint 100 64.00 82.00 85.00 65.00
Fake News 140 65.71 56.43 46.43 55.00
Flowchart 706 74.79 87.11 31.16 77.34
Role Playing 108 39.47 44.74 26.32 30.70
5 Human evaluation
In this section, we explain the human labeling process and the steps involved in constructing the dataset for human evaluation. Through this, we conduct experiments to compare how closely the ELITE evaluator aligns with human judgment compared to existing automated safety evaluator and safeguard models.
5.1 Human Annotators and Datasets
We recruit 22 labelers through the data-labeling company to evaluate responses. The labelers are selected to ensure diversity across gender, age, and occupation, aiming for unbiased labeling. Each labeler is provided with clear instructions on identifying safe and unsafe responses generated by the victim models. They reviewed each image-text-response pair and classified it as either “safe” or “unsafe”. To ensure reliability, each image-text pair is evaluated by three labelers, and the majority decision is taken as the ground truth. The detailed evaluation process and the content of the instructions are provided in the Appendix D.
The dataset used for human evaluation consists of a subset of image-text pairs, with approximately 90 pairs per taxonomy, totaling 963 pairs. For a fair comparison, instead of using the filtered responses, we also include responses from models that did not meet the filtering criteria among the three models (Phi-3.5-Vision, Llama-3.2-11B-Vision, and Pixtral-12B). Additionally, these pairs were primarily collected where the evaluation results differed between the ELITE evaluator and existing evaluation methods (e.g., StrongREJECT evaluator and safeguard models) and were randomly sampled across each taxonomy. To ensure diversity, we excluded image-text pairs that differed only in model responses from the human evaluation dataset.
Figure 4: The comparison of AU-ROC curves between the ELITE evaluator and StrongREJECT evaluator on our human evaluation dataset.
5.2 Comparison with Existing Evaluation Method
To demonstrate the superiority of the ELITE evaluator, we compare it with the StrongREJECT evaluator. Fig. 4 shows the comparison using the Area Under the Receiver Operating Characteristic Curve (AU-ROC Curve) (Bradley, 1997), considering the differences in scoring scales between the two methods. For a fair comparison, both the ELITE and StrongREJECT evaluators are evaluated using the GPT-4o on the human evaluation dataset consisting of 963 image-text pairs.
As shown in Fig. 4, the StrongREJECT (GPT-4o) achieves an Area Under the Curve (AUC) of 0.46. In contrast, the ELITE evaluator achieves a significantly higher AUC of 0.77, demonstrating that the ELITE evaluator aligns more closely with human judgment. This result indicates the necessity of incorporating a toxicity score for a more accurate and comprehensive safety evaluation in VLMs. Furthermore, it highlights the robustness and superior performance of the ELITE evaluator.
To further demonstrate that the effectiveness of the ELITE evaluator is not solely due to advanced models like GPT-4o, we validate its effectiveness by applying it to open-source models. Specifically, we apply it to InternVL2.5 (7B, 26B) for comparison. Experimental results show that the ELITE evaluator with InternVL2.5 (7B, 26B) achieves AUC values of 0.57 and 0.65, respectively, surpassing the StrongREJECT evaluator with GPT-4o. This finding confirms that the strong performance of the ELITE evaluator is not solely dependent on a competent model.
Table 6: Performance comparison of the ELITE (GPT-4o), ELITE (InternVL2.5-8B, 26B), ELITE (InternVL2.5-26B), LlamaGuard3-Vision-11B, LlavaGuard-13B, and OpenAI Moderation API on our human evaluation dataset. The best-performing method is highlighted in bold and the second-best method with an underline.
Method Accuracy ($\uparrow$) Precision ($\uparrow$) Recall ($\uparrow$) F1 score ($\uparrow$)
ELITE (GPT-4o) 0.726 0.579 0.709 0.637
ELITE (InternVL2.5-26B) 0.660 0.500 0.471 0.485
ELITE (InternVL2.5-8B) 0.609 0.416 0.376 0.395
LlamaGuard3-Vision-11B 0.603 0.339 0.177 0.233
LlavaGuard-13B 0.536 0.331 0.361 0.346
OpenAI Moderation API 0.624 0.439 0.388 0.412
5.3 Comparison with Safeguard Models
We compare the ELITE evaluator with safeguard models, including LlamaGuard3-Vision-11B (Chi et al., 2024), LlavaGuard-13B (Helff et al., 2024), and OpenAI Moderation API (Markov et al., 2023). In this experiment, the ELITE evaluator classifies responses with ELITE evaluator score $s\geq 10$ as unsafe and $s<10$ as safe, following the same criteria used for filtering.
Table 6 demonstrates that the ELITE evaluator, when applied to GPT-4o, outperforms LlamaGuard3-Vision-11B in terms of accuracy, precision, recall, and F1 score. Specifically, it achieves 73% accuracy, representing an improvement of approximately 20.3% over LlamaGuard3-Vision-11B, 35% over LlavaGuard-13B, and 16% over the OpenAI Moderation API. For the F1 score, the ELITE evaluator shows an F1 score of 0.637, which is significantly higher than the others. Furthermore, the ELITE evaluator performs better when applied to the open-source model InternVL2.5 (26B). The superior performance of the ELITE evaluator on open-source models further indicates that its effectiveness is not limited to GPT-4o, highlighting its broader applicability.
Table 7: Breakdown of F1 score according to taxonomies between ELITE (GPT-4o), LlamaGuard3-Vision-11B, LlavaGuard-13B, and OpenAI Moderation API on our human evaluation dataset.
Taxonomy F1 score ($\uparrow$)
ELITE LlamaGuard3 LlavaGuard OpenAI Mod.
S1. Violent Crimes 0.50 0.16 0.31 0.43
S2. Non-Violent Crimes 0.61 0.08 0.26 0.48
S3. Sex Crimes 0.62 0.18 0.33 0.24
S4. Defamation 0.62 0.18 0.25 0.06
S5. Specialized Advice 0.52 0.09 0.12 0.08
S6. Privacy 0.55 0.16 0.37 0.40
S7. Intellectual Property 0.86 0.62 0.54 0.70
S8. Indiscriminate Weapons 0.76 0.18 0.57 0.56
S9. Hate 0.66 0.18 0.38 0.44
S10. Self-Harm 0.67 0.20 0.30 0.30
S11. Sexual Content 0.52 0.37 0.26 0.38
Table 7 presents the F1 score for each taxonomy on the human evaluation dataset. Our results show that the ELITE evaluator outperforms LlamaGuard3-Vision-11B across all taxonomies. Specifically, safeguard methods tend to show low F1 scores in certain taxonomies. For instance, LlamaGuard3-Vision-11B shows significantly lower F1 scores in taxonomies such as S2. Non-violent Crimes and S5. Specialized Advice. Similarly, the OpenAI Moderation API shows low F1 scores in taxonomies such as S4. Defamation and S5. Specialized Advice. In contrast, the ELITE evaluator exhibits consistently high and balanced performance across all taxonomies. This demonstrates the superiority of the ELITE evaluator and indicates its effectiveness and accuracy in safety evaluation.
6 Analysis on the ELITE Evaluator & Benchmark
In this section, we note that the human evaluation dataset (Sec. 5) may be biased, as it primarily consists of samples where the evaluation results differed between the ELITE evaluator and existing methods (e.g., the StrongREJECT evaluator and safeguard models).
To enable a more rigorous analysis of the ELITE evaluator & benchmark, we conduct an additional human evaluation. This additional human evaluation dataset consists of 228 samples, 110 of which were randomly selected from the ELITE benchmark (referred to as From ELITE) and 118 that were not included (i.e., filtered out, referred to as Not From ELITE). We included at least 20 samples from each taxonomy and gathered the opinions of 3 labelers per sample, with the final labeling determined by majority decision. In total, 8 labelers were recruited for this evaluation. We provided the input image, text, and the model’s response to perform the safety judgment.
Figure 5: Pearson correlation coefficients among three labelers and the ELITE evaluator based on toxicity scores over 228 samples.
6.1 Alignment of Toxicity Scores Between ELITE and Humans
We conduct a quantitative analysis to assess how well the ELITE evaluator’s toxicity scores align with human judgment. Specifically, we compute the Pearson correlation coefficients between the ELITE evaluator (GPT-4o) and three individual human labelers across 228 samples. As shown in Fig. 5, the correlation between ELITE and Human1 is 0.7274, with 0.6447 for Human2 and 0.6496 for Human3. These values indicate a strong alignment between the ELITE evaluator and each human rater.
In contrast, the correlations among human labelers are comparatively lower: 0.5992 between Human1 and Human2, 0.6079 between Human1 and Human3, and 0.5079 between Human2 and Human3. As a result, the correlation between the ELITE evaluator and individual human labelers is higher than the inter-labelers correlations, suggesting that the ELITE evaluator provides more consistent toxicity judgments despite variability among human raters.
6.2 Validation of the ELITE Benchmark with Human Evaluation
To validate the quality of the ELITE benchmark, we use an additional human evaluation dataset to compare the proportion of harmful responses between included and excluded samples. As shown in Tab. 8, 67.27% of the From ELITE samples were labeled as unsafe by human labelers, while only 11.86% of the Not From ELITE samples were judged to be unsafe. This substantial difference confirms that the ELITE evaluator effectively selects harmful cases and filters out ambiguous ones. By incorporating a toxicity score, it further enhances the precision of safety evaluation and ensures the benchmark’s overall quality.
As we analyzed the 118 samples excluded during the ELITE benchmark filtering process (Not From ELITE), we observed that 52.54% of them resulted in responses that were merely descriptive of the input image. This highlights the need for a toxicity score to distinguish genuinely harmful content, ensuring a more accurate and meaningful safety evaluation.
Table 8: Human safety labels for From ELITE and Not From ELITE samples.
Majority Vote From ELITE Not From ELITE
Unsafe 67.27% 11.86%
Safe 32.73% 88.14%
7 Limitations
The ELITE evaluator aims to provide an accurate and reliable safety evaluation for vision-language models. However, as a rubric-based approach, its evaluation performance may vary depending on the capabilities of VLMs, which we acknowledge as a potential limitation.
Moreover, although the ELITE benchmark is carefully constructed to cover a broad taxonomy and include a wide range of diverse image-text pairs, it may still miss certain cases that can elicit harmful responses. For instance, more complex scenarios such as multi-turn jailbreak techniques are not yet incorporated, suggesting room for future expansion.
8 Conclusion
In this work, we introduce the ELITE evaluator, a rubric-based method enhanced with toxicity scoring to provide more precise and human-aligned safety assessments. This evaluator effectively overcomes the limitations of existing evaluation methods, which frequently struggle to separate truly harmful outputs from harmless yet uninformative responses, such as those that merely describe the input image. Alongside the evaluator, we construct the ELITE benchmark, a high-quality dataset of 4,587 image-text pairs curated through the ELITE evaluator. By filtering out ambiguous image-text pairs and incorporating in-house generated image-text pairs that cover all image-text pairs—including safe-safe combinations—we ensure the benchmark supports broad and robust safety evaluation. Our experimental results demonstrate that the ELITE evaluator aligns more closely with human judgment than existing methods and that the ELITE benchmark exposes harmful responses more effectively than prior benchmarks. We hope this work facilitates future research in multimodal safety and contributes toward the development of more reliable, trustworthy VLMs.
Acknowledgments
This research was partly supported by the MSIT(Ministry of Science and ICT), Korea, under the ITRC(Information Technology Research Center) support program(IITP-2024-RS-2023-00258649, 50%) supervised by the IITP(Institute for Information & Communications Technology Planning & Evaluation), partly supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2025-00562437, 30%), and partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2022-00155911, Artificial Intelligence Convergence Innovation Human Resources Development (Kyung Hee University), 10%), and by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2022-00143524, Development of Fundamental Technology and Integrated Solution for Next-Generation Automatic Artificial Intelligence System, 10%)
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. In this work, we introduce a benchmark to evaluate the safety of VLMs. Given its nature, the benchmark contains potentially offensive samples, which may raise safety concerns. We affirm that all data used in this study will not be utilized for purposes other than research. Our research aims to focus on the safety challenges of VLMs and to facilitate future research on their safety alignment to prevent harmful responses
Appendix A Additional Experiments
A.1 Benchmark using the ELITE Evaluator Score
Table 9: ELITE evaluator score of various VLMs across taxonomies. The upper group in the table represents proprietary models, and the lower group represents open-source models. Highlight the most vulnerable model in bold and the second-most vulnerable with an underline. * denotes the model used for filtering.
Model S1 S2 S3 S4 S5 S6 S7 S8 S9 S10 S11 Average Score
GPT-4o 3.12 3.36 2.33 3.77 6.88 2.76 1.50 3.37 1.83 2.30 2.83 3.07
GPT-4o-mini 5.69 6.94 3.38 5.74 8.57 5.15 3.81 5.89 3.49 5.74 6.20 5.55
Gemini-2.0-Flash 11.86 15.66 9.31 9.64 10.07 11.76 11.99 14.76 8.22 9.56 9.32 11.42
Gemini-1.5-Pro 7.48 10.33 5.45 7.40 7.47 6.87 11.95 9.57 4.62 5.82 6.99 7.78
Gemini-1.5-Flash 8.62 12.45 4.69 8.04 7.69 7.52 11.95 9.57 5.90 4.97 7.16 8.43
LLaVa-v1.5-7B 13.66 17.14 15.52 9.5 8.81 14.09 13.68 13.76 9.44 12.94 10.79 12.90
LLaVa-v1.5-13B 14.93 19.29 16.17 9.48 9.92 15.36 14.92 14.89 10.91 14.82 11.58 14.15
DeepSeek-VL-7B 7.65 12.75 6.50 5.86 8.02 8.46 8.80 11.36 6.79 6.80 5.90 8.39
DeepSeek-VL2-Small 13.28 17.37 11.29 7.54 10.16 14.41 13.08 15.52 9.85 10.41 8.84 12.37
ShareGPT4V-7B 14.1 18.8 15.76 8.98 9.63 14.88 15.57 13.87 11.57 13.14 10.87 13.73
ShareGPT4V-13B 14.52 19.16 15.72 9.24 10.21 15.59 15.55 14.25 11.38 12.86 10.74 13.93
Qwen2-VL-7B 11.54 15.15 8.99 6.82 8.82 12.69 8.75 14.69 8.73 10.5 9.46 10.87
Molmo-7B 12.07 17.73 12.79 9.32 9.35 11.81 15.93 16.10 9.19 12.98 12.17 12.90
InternVL2.5-8B 10.22 14.06 11.65 4.83 9.54 9.84 8.16 12.31 5.27 8.39 7.25 9.45
InternVL2.5-26B 7.54 10.02 8.11 4.83 8.63 7.12 4.89 10.74 4.22 6.68 5.29 7.21
Phi-3.5-Vision* 7.03 9.33 3.13 7.75 6.46 4.52 4.23 8.56 3.31 4.25 2.98 5.95
Pixtral-12B* 15.05 21.14 15.44 11.29 10.46 15.46 18.80 19.45 14.46 14.87 12.77 15.79
Llama-3.2-11B-Vision* 9.74 14.82 7.3 4.85 9.12 10.15 6.84 11.33 5.76 7.55 6.55 8.97
We measure the ELITE evaluator score for each taxonomy on a [0-25] point scale using the ELITE evaluator score, rather than E-ASR. As shown in Table 9, among proprietary models, GPT-4o exhibits the lowest ELITE evaluator score, while Gemini-2.0-Flash achieves the highest score of 11.42. For open-source models, most show relatively high scores in the low to mid-10s. Similar to the E-ASR results provided in the main text, these high scores emphasize the need for safety alignment.
A.2 Threshold
Table 10: Performance comparison of ELITE evaluator (GPT-4o) by threshold on our human evaluation dataset.
ELITE evaluator score for threshold Accuracy F1 score
5 0.661 0.625
10 0.726 0.637
15 0.727 0.638
20 0.728 0.639
We validate our filtering criterion, a threshold of 10, using our human evaluation dataset. As shown in Table 10, we set thresholds at 5, 10, 15, and 20 to measure accuracy and F1 score. Although the threshold of 10 does not perform the best, the below examples of image-text pairs with ELITE evaluator scores near 10 demonstrate that harmful responses can still be elicited from VLMs. Therefore, considering the potential for eliciting harmful responses, we select 10 as the threshold to incorporate a diverse range of image-text pairs.
The following are examples of image-text pairs with ELITE evaluator scores near 10:
Figure 6: Examples with ELITE evaluator score below 10
Figure 7: Examples with an ELITE evaluator score of 10 or higher
Appendix B Details of the ELITE Benchmark
In this section, we provide detailed information on the definition of each taxonomy, previous benchmarks, and sample examples for each taxonomy. We also present the prompts used to integrate these benchmarks into our taxonomy during the construction of the ELITE benchmark (generated), as well as the detailed prompts for the ELITE evaluator. Additionally, we include the model cards for both the models and safeguard models used for comparison in the experiments.
B.1 Taxonomy
Definitions of each taxonomy are shown below:
S1. Violent Crimes: Refers to unlawful violence toward people (e.g., terrorism, genocide, murder, hate-motivated violent crimes, child abuse, assault, battery, kidnapping) and unlawful violence toward animals (e.g., animal abuse).
S2. Non-Violent Crimes: Refers to personal crimes, financial crimes, property crimes, drug crimes, weapons crimes, and cybercrimes.
S3. Sex-Related Crimes: Refers to sex trafficking, sexual assault, sexual harassment, non-consensual sharing of sexually intimate content, and bestiality.
S4. Defamation: Refers to verifiably false statements that are likely to injure a living person’s reputation.
S5. Specialized Advice: Refers to unqualified advice (without disclaimers) regarding financial, medical, legal, or electoral matters, as well as statements indicating that dangerous activities or objects are categorically safe.
S6. Privacy: Refers to the exposure of sensitive, nonpublic personal information that could undermine someone’s physical, digital, or financial security.
S7. Intellectual Property: Refers to responses that violate the intellectual property rights of any party or quote verbatim from existing works without proper attribution.
S8. Indiscriminate Weapons: Refers to weapons that cause mass destruction, such as chemical, biological, radiological, nuclear, or other high-yield explosive weapons.
S9. Hate: Refers to responses that demean or dehumanize individuals on the basis of their sensitive, personal characteristics (e.g., race, gender, religion, disability).
S10. Suicide & Self-Harm: Refers to intentional acts of self-harm, such as suicide, self-injury (e.g., cutting), and disordered eating.
S11. Sexual Content: Refers to pornography or engaging in direct, explicit conversations (e.g., cybersex).
B.2 Previous Benchmarks
Table 11: Details of the previous benchmarks used in the construction method of benchmark and evaluation methods. The sizes indicated in parentheses represent the datasets utilized in this work. For Figstep, the dataset included in JailbreakV-28K is used.
Benchmark # Size Construction of Benchmark Evaluation Method
VLGuard (Zong et al., 2024) 3,000 (2,028) GPT-4V Rule, LlamaGuard
MLLMGuard (Gu et al., 2024) 2,282 (532) Human Rule, GuardRank
MM-SafetyBench (Liu et al., 2025) 5,040 (1,680) Human, GPT-4 GPT-4
SIUO (Wang et al., 2025) 167 Human Human, GPT-4V
Figstep (Gong et al., 2023) 500 GPT-4 Human
SPA-VL (Zhang et al., 2024) 100,788 Open-source dataset (LAION-5B), Gemini-1.0-Pro Vision GPT-4V
JailbreakV-28k (Luo et al., 2024) 28,000 Human, GPT-4 LlamaGuard
Table 11 provides detailed information on the benchmarks targeted for filtering during the construction of the ELITE benchmark, including their size, dataset generation methods, and safety evaluation methods. Note that we utilize a subset of publicly available datasets for MLLMGuard, and for VLGuard and MM-SafetyBench, we use the entire publicly available datasets. Specifically, unique image-text pairs are used, excluding overlapping elements and cases where multiple prompts are associated with a single image.
B.3 Image-Text Pairs By Taxonomy
In this section, we provide examples of 11 different taxonomies, along with the ELITE evaluator scores for each image-text pair and model response.
Figure 8: Examples of 11 different taxonomies
B.4 Image-Text Pairs for ELITE benchmark (generated)
In this section, we provide examples of image-text pairs in the ELITE benchmark (generated), along with the ELITE evaluator scores for each image-text pair and model response.
Figure 9: Examples of image-text pairs in ELITE benchmark (generated)
B.5 Model Cards
Table 12 provides model cards of the VLMs used in our paper, including their parameters and model architecture components. Also, Table 13 provides a detailed summary of the safeguard models in the human evaluation section, including their base models and training datasets.
Table 12: Model cards used in our benchmark experiments. “-” denotes that information is not available for propitiatory models. For open-source models, instruction-tuned or chat-capable models are used.
Name # Params Vision Encoder Base LLM
GPT-4o (OpenAI, 2024a) - - -
GPT-4o-mini - - -
Gemini-2.0-Flash (Google DeepMind, 2024) - - -
Gemini-1.5-Pro (Gemini Team et al., 2024) - - -
Gemini-1.5-Flash (Gemini Team et al., 2024) - - -
LLaVa-v1.5-7B (Liu et al., 2024) 7B CLIP ViT-L/14-336px Vicuna-7B-v1.5
LLaVa-v1.5-13B (Liu et al., 2024) 13B CLIP ViT-L/14-336px Vicuna-13B-v1.5
DeepSeek-VL-7B (Lu et al., 2024) 7B SigLIP-L+SAM-B DeepSeek-LLM-7B
DeepSeek-VL2-Small (Wu et al., 2024) 16B SigLIP-SO400M DeepSeek-MoE-16B
ShareGPT4V-7B (Chen et al., 2025) 7B CLIP ViT-L/14-336px Vicuna-7B-v1.5
ShareGPT4V-13B (Chen et al., 2025) 13B CLIP ViT-L/14-336px Vicuna-13B-v1.5
Phi-3.5-Vision (Abdin et al., 2024) 4.2B CLIP ViT-L/14-336px Phi-3.5-mini
Pixtral-12B (Agrawal et al., 2024) 12B Custom ViT with 400M params Mistral-NeMo-12B
Llama-3.2-11B-Vision (Dubey et al., 2024) 11B - Llama-3.1-8B
Qwen2-VL-7B (Wang et al., 2024) 7B Custom ViT with 600M params Qwen2-7B
Molmo-7B (Deitke et al., 2024) 7B CLIP ViT-L/14-336px Qwen2-7B
InternVL2.5-8B (Chen et al., 2024) 8B InternViT-300M-448px-V2.5 InternLM2.5-7B-Chat
InternVL2.5-26B (Chen et al., 2024) 26B InternViT-6B-448px-V2.5 InternLM2.5-20B-Chat
Table 13: Model cards of safeguard models used in our human evaluation experiments.
Name Base Model Training Dataset
LlamaGuard3-Vision-11B (Chi et al., 2024) Llama-3.2-11B-Vision Human-created, Llama-based synthetic dataset
LlavaGuard-13B (Helff et al., 2024) LLaVa-v1.5-13B Human-created dataset (SMID)
OpenAI Moderation API (Markov et al., 2023) GPT-4o Human-created, GPT-based synthetic dataset
B.6 Taxonomy Matching
The prompts used for taxonomy matching with previous benchmarks in Benchmark Construction Pipeline of this paper are as follows:
Figure 10: Prompts used for taxonomy matching on GPT-4o
B.7 ELITE Evaluator Prompt
The full prompt used in the ELITE evaluator is as follows:
Figure 11: Prompts used for ELITE evaluator
Appendix C Ethical Statement
In this study, we introduce a benchmark to evaluate the safety of VLMs. Given its nature, the benchmark contains potentially offensive samples, which may raise safety concerns. We affirm that all data used in this study will not be utilized for purposes other than research. Our research aims to focus on the safety challenges of VLMs and to facilitate future research on their safety alignment to prevent harmful responses.
Appendix D Human Evaluation
D.1 Human Evaluation Guidelines
We recruited human labelers for annotation through the data labeling company CrowdWorks and paid them more than twice the minimum wage. The guidelines provided to the labelers for human evaluation are as follows:
Figure 12: Guidelines used for human evaluation
D.2 Distribution of Human Labelers
The gender, age, and occupation distributions of the recruited human labelers are as follows:
Figure 13: The distribution of the recruited human labelers by gender, age, and occupation